# A Comprehensive Survey on World Models for Embodied AI

### v1

@@ 3.1.2 Global Difference Prediction
On the representation side, TOKEN [95] tokenizes scenes into object-level tokens, aligning world representations with reasoning and leveraging LLMs to predict full future trajectories for long-tail scenarios. GeoDrive [170] extracts 3D representations, renders trajectory-conditioned views, and edits the position of vehicles to guide DiT in producing editable generations. For control, FLARE [171] aligns diffusion policies with latent future representations, avoiding pixel-space video generation and learning effectively from action-free videos. In a related vein, LaDi-WM [172] predicts future states through interactive diffusion in a latent space aligned with visual foundation models, integrating geometric and semantic features while iteratively refining the diffusion policy to improve performance and generalization. villa-X [76] and VidMan [55] both couple diffusion-based models with IDM for control. villa-X infers latent actions, aligns them with ego-centric forward dynamics, and maps them via joint diffusion, while VidMan adapts a pretrained video-diffusion model into an IDM using a self-attention adapter for accurate action prediction.
@@
Diffusion-based world models are commonly used for parallel generation. EmbodiedDreamer [173] couples differentiable physics with video diffusion to render photorealistic and physically consistent futures. TesserAct [78] reconstructs 4D spatiotemporal consistent scenes by jointly generating RGB, depth, and normal videos for IDM-based action learning. DFIT-OccWorld [174] reformulates prediction as decoupled voxel warping and adopts image-assisted single-stage training for reliable and efficient dynamic scene modeling. For instruction-conditioned control, RoboDreamer [49] decomposes instructions into low-level primitives that steer video diffusion, synthesizing novel compositional scenes beyond the training distribution while grounding execution via an IDM, and ManipDreamer [175] extends this design with an action-tree prior together with depth and semantic guidance to improve instruction following and temporal consistency.
On the planning side, 3DFlowAction [176] employs a pretrained 3D optical flow world model to treat future motion as a unified action cue, enabling label-free and cross-robot manipulation through closed-loop optimization. Imagine-2-Drive [177] integrates video diffusion with a multimodal diffusion policy to accelerate policy learning. Drive-WM [88] uses multi-view diffusion with image-based rewards to select safer trajectories, while World4Drive [178] leverages vision-based priors to construct intent-aware world models that support self-supervised multi-intent imagination. COMBO [179] composes multi-agent actions with diffusion, leverages VLMs to infer purposes, and integrates tree search for online cooperative planning.
3.2 General-Purpose World Models
3.2.1 Sequential Simulation and Inference
@@
Some general world models increasingly pretrain on unlabeled video and use tokenized latent space for robust forecasting and generation. iVideoGPT [56] was pretrained on large-scale interaction videos for action-free forecasting and later adapted to downstream control. Genie [50] learned discrete latent actions and spatiotemporal tokens, enabling user-controllable interactive environments via autoregressive dynamics. RoboScape [180] jointly learned video generation with temporal depth and keypoint dynamics to improve physical realism. PACT [181] tokenized multimodal perception and action and trained a causal Transformer to obtain a unified representation for diverse tasks, and DINO-world [182] learns generalizable dynamics by predicting the temporal evolution of DINOv2 features from large-scale unlabeled video corpora. Building on language priors, EVA [71] introduced a Reflection-of-Generation (RoG) policy that used a VLM for iterative self-correction, strengthening long-horizon anticipation. In the same vein, Owl-1 [183] employs a VLM to forecast world dynamics conditioned on current states and generated fragments, explicitly guiding the subsequent fragments and enabling coherent long-horizon video synthesis, while World4Omni [68] employs a reflective world model, where a VLM refines subgoal images from an image generator, and integrates them with pretrained modules for zero-shot robotic manipulation.
@@
Recent work adapts video diffusion models into controllable world models that autoregressively imagine future scenes. AdaWorld [72] introduced an action-aware pretraining scheme that extracted self-supervised latent actions between adjacent frames to condition diffusion, enabling efficient transfer with minimal interaction. Vid2World [184] adapts pretrained video diffusion models into autoregressive interactive world models via causalization and a causal action-guidance mechanism. GenAD [90] employs a two-stage strategy to adapt diffusion into a general video-prediction model conditioned on text and actions, enabling large-scale driving simulation and planning. Pandora [47] uses an instruction-tuned LLM to autoregressively steer a separate video-diffusion generator for explicit, goal-directed control, while Yume [75] quantized camera motions into text tokens to guide a masked Video DiT, enabling autoregressive synthesis of dynamic 3D exploratory worlds.
@@
To maintain geometric fidelity and long-horizon stability, recent methods couple explicit 3D priors with temporal-consistency modules in diffusion-based world models. At the geometric level, Geometry Forcing [185] aligns latent features with a geometric foundation model to inject explicit 3D priors, improving geometric consistency, while DeepVerse [64] integrates visual and geometric prediction targets and introduces a geometry-aware memory to sustain consistent long-horizon generation. For temporal stability, VRAG [186] proposed a Video Retrieval-Augmented Generation (RAG) framework that retrieves historical frames conditioned on a global state to stabilize autoregressive rollouts, StateSpaceDiffuser [63] combines Mamba with diffusion to alleviate long-term memory loss and content drift under short context windows, and InfinityDrive [187] injects memory and adopts an adaptive loss within DiT, producing minute-scale driving videos with high fidelity, temporal consistency, and diverse content. Complementing these designs, LongDWM [188] mitigates error accumulation in long-horizon video generation through distillation—where a fine-grained DiT learns continuous motion to guide a coarse model, whereas MiLA [189] adopts a coarse-to-fine strategy that predicts sparse anchor frames and refines them during interpolation to improve temporal consistency and long-term fidelity. Finally, for dynamics and conditioning, Orbis [190] employs a continuous-space flow-matching formulation that demonstrates greater robustness than discrete-token schemes for long-horizon rollouts, and DriVerse [109] leverages multimodal trajectory prompts with latent motion alignment to synthesize long-horizon driving videos from a single image and navigation trajectories.
@@
Sequential world models increasingly act as learned simulators, providing action-conditioned rollouts for policy evaluation and training. WorldGym [191] and WorldEval [192] generate action-conditioned rollouts and use VLM-based critics for evaluation, while WorldEval further leverages latent action representations to drive a DiT-based synthesizer. RLVR-World [62] fine-tunes world models with RL with Verifiable Rewards (RLVR), using explicit metrics to close the pretraining–task objective gap. For safety risk prediction, Guan et al. [193] presented a framework for autonomous driving accident prediction that augments data with a domain-informed world model and enhances spatiotemporal reasoning using graph and temporal convolutions.
@@
Beyond diffusion, sequence models broadened the capacity for long-range consistency. Po et al. [194] integrated block-wise state space models for long-term memory with local attention for short-term coherence, enabling video generation with sustained memory and consistent dynamics. S2-SSM [61] employs a Mamba layer to model the independent evolution of object slots and a sparsity-regularized cross-attention mechanism to capture their causal interactions, enabling causal reasoning over the environment.
@@
Building on this paradigm, structured-grid and physics-informed methods encode geometry and dynamics for controllable rollouts. PhyDNet [195] disentangles physical priors expressed as partial differential equations from visual factors, improving prediction. ViDAR [89] unifies semantics, geometry, and dynamics through a pre-training task of point cloud forecasting and a latent rendering operator, enabling a scalable foundation for downstream autonomous driving tasks. FOLIAGE [196] models dynamics with an Accretive Graph Network and a Transformer-based predictor, executing rollouts on simulated data. Complementing these grid and physics lines, MindJourney [73] couples VLMs with a controllable world model to render egocentric rollouts along planned camera trajectories, enabling multi-view reasoning.
@@
Building on grid-based representations, diffusion-based forecasting has become dominant for stable long-horizon generation. Within grid-centric predictors, DOME [94] encodes observations into a continuous latent space and applies a spatiotemporal DiT for scene forecasting, Copilot4D [82] tokenizes point clouds and couples a spatiotemporal Transformer with discrete diffusion to improve fidelity and coherence, and LidarDM [100] generates layout-conditioned static scenes, composes them with dynamic objects, integrating LiDAR simulation to produce controllable videos. For long-video generation, Vista [96] adopts a two-stage large-scale training regime to produce controllable, high-fidelity, driving videos, and Delphi [86] enforces long-horizon multiview consistency through shared noise and feature alignment and a failure-driven framework to synthesize targeted data for improved planning. To strengthen long-horizon stability, GEM [102] achieves controllable ego-vision generation through large-scale training and fine-grained control over motion, dynamics, and posture, Zhou et al. [197] maintains a persistent RGB-D 3D memory map to guide subsequent frames, and STAGE [198] introduced hierarchical temporal feature transfer with multi-stage training.
@@
Decomposed Rendering Representation.  Scenes are decomposed into explicit primitives to synthesize view-consistent, simulatable trajectories over long horizons. Within this paradigm, GaussianWorld [103] models scene evolution as ego-motion, object dynamics, and newly observed regions, iteratively updating 3D Gaussian primitives to enable accurate and efficient dynamic perception. InfiniCube [22] introduced a hybrid pipeline that combines voxel-based generation, video synthesis, and dynamic Gaussian reconstruction, enabling large-scale dynamic 3D driving scenes conditioned on HDmaps, bounding boxes, and text. Complementarily, Wu et al. [199] augmented a video world model with long-term spatial memory grounded in reconstructed geometry and an episodic memory, which together condition sequential generation for long-range consistency.
3.2.2 Global Difference Prediction
@@
Within Joint-Embedding Predictive Architecture (JEPA) [200], V-JEPA [51] extends this architecture to video by predicting latent features of occluded spatiotemporal regions, learning generalizable representations for both appearance and motion without pixel reconstruction or contrastive learning. Building on this, V-JEPA 2 [14] scales pretraining to large-scale Internet videos with larger models and incorporates limited robot interaction data for post-training, transferring to robotic planning. AD-L-JEPA [201] adapts JEPA to BEV LiDAR, predicting masked embeddings in a self-supervised manner. Beyond JEPA-style prediction, WorldDreamer [45] frames world modeling as masked visual sequence prediction to learn physics and motion for diverse video generation and editing, while MaskGWM [104] combines diffusion with masked feature reconstruction and a dual-branch masking strategy to improve long-horizon consistency and generalization.
In parallel, diffusion-based methods have become central to global-difference modeling. Sora [13] represents video as unified spacetime patches and uses a DiT to generate long, coherent sequences at scale. ForeDiff [202] decouples conditioning from denoising by adding a deterministic predictive stream and employing a pretrained predictor to guide generation, improving accuracy and consistency. For domain-specific synthesis, AirScape [203] introduces an aerial video–intention dataset, applies supervised fine-tuning for controllability, and leverages VLMs to impose spatiotemporal constraints; MarsGen [204] builds a multimodal Mars dataset from NASA’s sparse rover stereo imagery using geometric foundation models, then trains a controllable generator to produce visually realistic, geometry-consistent Martian videos. In clinical guidance, EchoWorld [205] proposes a motion-aware world model for echocardiography probe control, pretraining on region- and motion-outcome prediction and fine-tuning attention to fuse visual and motion cues for precise guidance.
@@
Recent work converges on unified scene understanding and future prediction. UniFuture [206] couples Dual Latent Sharing with multi-scale latent interaction to jointly model appearance and depth in future driving scenes, and HERMES [108] integrates multiview BEV features into an LLM with world queries that link scene understanding to future prediction within a single framework. BEVWorld [207] maps images and LiDAR into a compact BEV latent space through a unified tokenizer and applies a latent BEV diffusion model for synchronized multimodal forecasting. Progress in grid and occupancy forecasting includes differentiable raycasting with a proxy reformulation for sensor agnostic motion learning by Khurana et al. [208, 209] and LiDAR to range images 3D spatiotemporal convolutions by Mersch et al. [210]. Cam4DOcc [211] established the first vision-only benchmark with an E2E 3D CNN baseline, and Liu et al. [212] enhanced cross-task transfer through high-ratio compression and latent flow matching.
On the generative front, tokenized 4D representations enable controllable scene synthesis. OccSora [85] uses a 4D tokenizer to derive compact representations for trajectory conditioned diffusion, and DynamicCity [99] encodes 4D occupancy as HexPlanes representations with a VAE and employs a conditional DiT for high fidelity controllable dynamics. Fidelity and consistency improve through decoupling ego-motion with scene evolution in COME [213], physics-informed constraints in DrivePhysica [214], cross-view point map alignment in Liu et al. [215], and photometric warping-based supervision in PosePilot [216]. For controllable conditioning, DriveDreamer 2 [97] translates prompts into agent trajectories and HDMaps for customizable video generation, EOT-WM [217] encodes ego and surrounding trajectories as trajectory videos for trajectory consistent synthesis, and ORV [65] uses 4D semantic occupancy sequences to guide action conditioned video with S2R transfer. AETHER [77] unifies dynamic 4D reconstruction, action-conditioned video prediction, and vision-based planning under training on synthetic 4D data and achieves zero-shot generalization to real-world scenarios.
@@
A trend is combining video generation with Gaussian Splatting. DriveDreamer4D [105] exploits complex driving trajectories such as lane changes to guide video synthesis and optimize a 4DGS model, which enhances reconstruction fidelity and spatiotemporal coherence from novel viewpoints. ReconDreamer [106] introduces an online restoration module together with progressive data reuse to correct artifacts in Gaussian-rendered views and enables reliable reconstruction of large-scale trajectories. MagicDrive3D [84] generates multi-view street scenes conditioned on BEV maps, 3D boxes, and text, and further converts the outputs into full 3D environments through fault-tolerant GS. In contrast, implicit-field methods replace GS with continuous neural representations. UnO [17] leverages future point clouds to learn a NeRF-style 4D occupancy field, which allows annotation-free prediction and achieves strong transfer performance beyond supervised baselines in point-cloud forecasting.
4 Data Resources & Metrics
World models in embodied AI are required to address diverse tasks spanning manipulation, navigation, and autonomous driving, requiring heterogeneous resources and rigorous evaluation. Accordingly, we present Data Resources in §4.1 and Metrics in §4.2, focusing on the most widely adopted platforms and evaluation measures as a unified foundation for cross-domain assessment.
4.1 Data Resources

### v2

@@ III-A2 Global Difference Prediction
On the representation side, TOKEN [tian2025tokenize] tokenizes scenes into object-level tokens, aligning world representations with reasoning and leveraging LLMs to predict full future trajectories for long-tail scenarios. GeoDrive [chen2025geodrive] extracts 3D representations, renders trajectory-conditioned views, and edits the position of vehicles to guide DiT in producing editable generations. For control, FLARE [zheng2025flare] aligns diffusion policies with latent future representations, avoiding pixel-space video generation and learning effectively from action-free videos. In a related vein, LaDi-WM [huang2025ladi] predicts future states through interactive diffusion in a latent space aligned with visual foundation models, integrating geometric and semantic features while iteratively refining the diffusion policy to improve performance and generalization. villa-X [chen2025villa] and VidMan [wen2024vidman] both couple diffusion-based models with IDM for control. villa-X infers latent actions, aligns them with ego-centric forward dynamics, and maps them via joint diffusion, while VidMan adapts a pretrained video-diffusion model into an IDM using a self-attention adapter for accurate action prediction.
@@
Diffusion-based world models are commonly used for parallel generation. EmbodiedDreamer [wang2025embodiedreamer] couples differentiable physics with video diffusion to render photorealistic and physically consistent futures. TesserAct [zhen2025tesseract] reconstructs 4D spatiotemporal consistent scenes by jointly generating RGB, depth, and normal videos for IDM-based action learning. DFIT-OccWorld [zhang2024efficient] reformulates prediction as decoupled voxel warping and adopts image-assisted single-stage training for reliable and efficient dynamic scene modeling. For instruction-conditioned control, RoboDreamer [zhou2024robodreamer] decomposes instructions into low-level primitives that steer video diffusion, synthesizing novel compositional scenes beyond the training distribution while grounding execution via an IDM, and ManipDreamer [li2025manipdreamer] extends this design with an action-tree prior together with depth and semantic guidance to improve instruction following and temporal consistency.
On the planning side, 3DFlowAction [zhi20253dflowaction] employs a pretrained 3D optical flow world model to treat future motion as a unified action cue, enabling label-free and cross-robot manipulation through closed-loop optimization. Imagine-2-Drive [garg2025imagine] integrates video diffusion with a multimodal diffusion policy to accelerate policy learning. Drive-WM [wang2024driving] uses multi-view diffusion with image-based rewards to select safer trajectories, while World4Drive [zheng2025world4drive] leverages vision-based priors to construct intent-aware world models that support self-supervised multi-intent imagination. COMBO [zhang2025combo] composes multi-agent actions with diffusion, leverages VLMs to infer purposes, and integrates tree search for online cooperative planning.
III-B General-Purpose World Models
III-B1 Sequential Simulation and Inference
@@
Some general world models increasingly pretrain on unlabeled video and use tokenized latent space for robust forecasting and generation. iVideoGPT [wu2024ivideogpt] was pretrained on large-scale interaction videos for action-free forecasting and later adapted to downstream control. Genie [bruce2024genie] learned discrete latent actions and spatiotemporal tokens, enabling user-controllable interactive environments via autoregressive dynamics. RoboScape [shang2025roboscape] jointly learned video generation with temporal depth and keypoint dynamics to improve physical realism. PACT [bonatti2023pact] tokenized multimodal perception and action and trained a causal Transformer to obtain a unified representation for diverse tasks, and DINO-world [baldassarre2025back] learns generalizable dynamics by predicting the temporal evolution of DINOv2 features from large-scale unlabeled video corpora. Building on language priors, EVA [chi2025empowering] introduced a Reflection-of-Generation (RoG) policy that used a VLM for iterative self-correction, strengthening long-horizon anticipation. In the same vein, Owl-1 [huang2024owl] employs a VLM to forecast world dynamics conditioned on current states and generated fragments, explicitly guiding the subsequent fragments and enabling coherent long-horizon video synthesis, while World4Omni [chen2025world4omni] employs a reflective world model, where a VLM refines subgoal images from an image generator, and integrates them with pretrained modules for zero-shot robotic manipulation.
@@
Recent work adapts video diffusion models into controllable world models that autoregressively imagine future scenes. AdaWorld [gao2025adaworld] introduced an action-aware pretraining scheme that extracted self-supervised latent actions between adjacent frames to condition diffusion, enabling efficient transfer with minimal interaction. Vid2World [huang2025vid2world] adapts pretrained video diffusion models into autoregressive interactive world models via causalization and a causal action-guidance mechanism. GenAD [yang2024generalized] employs a two-stage strategy to adapt diffusion into a general video-prediction model conditioned on text and actions, enabling large-scale driving simulation and planning. Pandora [xiang2024pandora] uses an instruction-tuned LLM to autoregressively steer a separate video-diffusion generator for explicit, goal-directed control, while Yume [mao2025yume] quantized camera motions into text tokens to guide a masked Video DiT, enabling autoregressive synthesis of dynamic 3D exploratory worlds.
@@
To maintain geometric fidelity and long-horizon stability, recent methods couple explicit 3D priors with temporal-consistency modules in diffusion-based world models. At the geometric level, Geometry Forcing [wu2025geometry] aligns latent features with a geometric foundation model to inject explicit 3D priors, improving geometric consistency, while DeepVerse [chen2025deepverse] integrates visual and geometric prediction targets and introduces a geometry-aware memory to sustain consistent long-horizon generation. For temporal stability, VRAG [chen2025learning] proposed a Video Retrieval-Augmented Generation (RAG) framework that retrieves historical frames conditioned on a global state to stabilize autoregressive rollouts, StateSpaceDiffuser [savov2025statespacediffuser] combines Mamba with diffusion to alleviate long-term memory loss and content drift under short context windows, and InfinityDrive [guo2024infinitydrive] injects memory and adopts an adaptive loss within DiT, producing minute-scale driving videos with high fidelity, temporal consistency, and diverse content. Complementing these designs, LongDWM [wang2025longdwm] mitigates error accumulation in long-horizon video generation through distillation—where a fine-grained DiT learns continuous motion to guide a coarse model, whereas MiLA [wang2025mila] adopts a coarse-to-fine strategy that predicts sparse anchor frames and refines them during interpolation to improve temporal consistency and long-term fidelity. Finally, for dynamics and conditioning, Orbis [mousakhan2025orbis] employs a continuous-space flow-matching formulation that demonstrates greater robustness than discrete-token schemes for long-horizon rollouts, and DriVerse [li2025driverse] leverages multimodal trajectory prompts with latent motion alignment to synthesize long-horizon driving videos from a single image and navigation trajectories.
@@
Sequential world models increasingly act as learned simulators, providing action-conditioned rollouts for policy evaluation and training. WorldGym [quevedo2025worldgym] and WorldEval [li2025worldeval] generate action-conditioned rollouts and use VLM-based critics for evaluation, while WorldEval further leverages latent action representations to drive a DiT-based synthesizer. RLVR-World [wu2025rlvr] fine-tunes world models with RL with Verifiable Rewards (RLVR), using explicit metrics to close the pretraining–task objective gap. For safety risk prediction, Guan et al. [guan2025world] presented a framework for autonomous driving accident prediction that augments data with a domain-informed world model and enhances spatiotemporal reasoning using graph and temporal convolutions.
@@
Beyond diffusion, sequence models broadened the capacity for long-range consistency. Po et al. [po2025long] integrated block-wise state space models for long-term memory with local attention for short-term coherence, enabling video generation with sustained memory and consistent dynamics. S2-SSM [petri2025learning] employs a Mamba layer to model the independent evolution of object slots and a sparsity-regularized cross-attention mechanism to capture their causal interactions, enabling causal reasoning over the environment.
@@
Building on this paradigm, structured-grid and physics-informed methods encode geometry and dynamics for controllable rollouts. PhyDNet [guen2020disentangling] disentangles physical priors expressed as partial differential equations from visual factors, improving prediction. ViDAR [Yang2024visual] unifies semantics, geometry, and dynamics through a pre-training task of point cloud forecasting and a latent rendering operator, enabling a scalable foundation for downstream autonomous driving tasks. FOLIAGE [liu2025foliage] models dynamics with an Accretive Graph Network and a Transformer-based predictor, executing rollouts on simulated data. Complementing these grid and physics lines, MindJourney [yang2025mindjourney] couples VLMs with a controllable world model to render egocentric rollouts along planned camera trajectories, enabling multi-view reasoning.
@@
Building on grid-based representations, diffusion-based forecasting has become dominant for stable long-horizon generation. Within grid-centric predictors, DOME [gu2024dome] encodes observations into a continuous latent space and applies a spatiotemporal DiT for scene forecasting, Copilot4D [zhang2024copilot4d] tokenizes point clouds and couples a spatiotemporal Transformer with discrete diffusion to improve fidelity and coherence, and LidarDM [zyrianov2025lidardm] generates layout-conditioned static scenes, composes them with dynamic objects, integrating LiDAR simulation to produce controllable videos. For long-video generation, Vista [gao2024vista] adopts a two-stage large-scale training regime to produce controllable, high-fidelity, driving videos, and Delphi [ma2024unleashing] enforces long-horizon multiview consistency through shared noise and feature alignment and a failure-driven framework to synthesize targeted data for improved planning. To strengthen long-horizon stability, GEM [hassan2025gem] achieves controllable ego-vision generation through large-scale training and fine-grained control over motion, dynamics, and posture, Zhou et al. [zhou2025learning] maintains a persistent RGB-D 3D memory map to guide subsequent frames, and STAGE [wang2025stage] introduced hierarchical temporal feature transfer with multi-stage training.
@@
Decomposed Rendering Representation.  Scenes are decomposed into explicit primitives to synthesize view-consistent, simulatable trajectories over long horizons. Within this paradigm, GaussianWorld [zuo2025gaussianworld] models scene evolution as ego-motion, object dynamics, and newly observed regions, iteratively updating 3D Gaussian primitives to enable accurate and efficient dynamic perception. InfiniCube [lu2025infinicube] introduced a hybrid pipeline that combines voxel-based generation, video synthesis, and dynamic Gaussian reconstruction, enabling large-scale dynamic 3D driving scenes conditioned on HDmaps, bounding boxes, and text. Complementarily, Wu et al. [wu2025video] augmented a video world model with long-term spatial memory grounded in reconstructed geometry and an episodic memory, which together condition sequential generation for long-range consistency.
III-B2 Global Difference Prediction
@@
Within Joint-Embedding Predictive Architecture (JEPA) [LeCun2022Path], V-JEPA [bardes2024revisiting] extends this architecture to video by predicting latent features of occluded spatiotemporal regions, learning generalizable representations for both appearance and motion without pixel reconstruction or contrastive learning. Building on this, V-JEPA 2 [assran2025v] scales pretraining to large-scale Internet videos with larger models and incorporates limited robot interaction data for post-training, transferring to robotic planning. AD-L-JEPA [zhu2025ad] adapts JEPA to BEV LiDAR, predicting masked embeddings in a self-supervised manner. Beyond JEPA-style prediction, WorldDreamer [wang2024worlddreamer] frames world modeling as masked visual sequence prediction to learn physics and motion for diverse video generation and editing, while MaskGWM [ni2025maskgwm] combines diffusion with masked feature reconstruction and a dual-branch masking strategy to improve long-horizon consistency and generalization.
In parallel, diffusion-based methods have become central to global-difference modeling. Sora [brooks2024video] represents video as unified spacetime patches and uses a DiT to generate long, coherent sequences at scale. ForeDiff [zhang2025consistent] decouples conditioning from denoising by adding a deterministic predictive stream and employing a pretrained predictor to guide generation, improving accuracy and consistency. For domain-specific synthesis, AirScape [zhao2025airscape] introduces an aerial video–intention dataset, applies supervised fine-tuning for controllability, and leverages VLMs to impose spatiotemporal constraints; MarsGen [li2025martian] builds a multimodal Mars dataset from NASA’s sparse rover stereo imagery using geometric foundation models, then trains a controllable generator to produce visually realistic, geometry-consistent Martian videos. In clinical guidance, EchoWorld [yue2025echoworld] proposes a motion-aware world model for echocardiography probe control, pretraining on region- and motion-outcome prediction and fine-tuning attention to fuse visual and motion cues for precise guidance.
@@
Recent work converges on unified scene understanding and future prediction. UniFuture [liang2025seeing] couples Dual Latent Sharing with multi-scale latent interaction to jointly model appearance and depth in future driving scenes, and HERMES [zhou2025hermes] integrates multiview BEV features into an LLM with world queries that link scene understanding to future prediction within a single framework. BEVWorld [zhang2024bevworld] maps images and LiDAR into a compact BEV latent space through a unified tokenizer and applies a latent BEV diffusion model for synchronized multimodal forecasting. Progress in grid and occupancy forecasting includes differentiable raycasting with a proxy reformulation for sensor agnostic motion learning by Khurana et al. [khurana2022differentiable, khurana2023point] and LiDAR to range images 3D spatiotemporal convolutions by Mersch et al. [mersch2022self]. Cam4DOcc [ma2024cam4docc] established the first vision-only benchmark with an E2E 3D CNN baseline, and Liu et al. [liu2025towards] enhanced cross-task transfer through high-ratio compression and latent flow matching.
On the generative front, tokenized 4D representations enable controllable scene synthesis. OccSora [wang2024occsora] uses a 4D tokenizer to derive compact representations for trajectory conditioned diffusion, and DynamicCity [bian2025dynamiccity] encodes 4D occupancy as HexPlanes representations with a VAE and employs a conditional DiT for high fidelity controllable dynamics. Fidelity and consistency improve through decoupling ego-motion with scene evolution in COME [shi2025come], physics-informed constraints in DrivePhysica [yang2024physical], cross-view point map alignment in Liu et al. [liu2025geometry], and photometric warping-based supervision in PosePilot [jin2025posepilot]. For controllable conditioning, DriveDreamer 2 [zhao2025drivedreamer] translates prompts into agent trajectories and HDMaps for customizable video generation, EOT-WM [zhu2025other] encodes ego and surrounding trajectories as trajectory videos for trajectory consistent synthesis, and ORV [yang2025orv] uses 4D semantic occupancy sequences to guide action conditioned video with S2R transfer. AETHER [team2025aether] unifies dynamic 4D reconstruction, action-conditioned video prediction, and vision-based planning under training on synthetic 4D data and achieves zero-shot generalization to real-world scenarios.
@@
A trend is combining video generation with Gaussian Splatting. DriveDreamer4D [zhao2025drivedreamer4d] exploits complex driving trajectories such as lane changes to guide video synthesis and optimize a 4DGS model, which enhances reconstruction fidelity and spatiotemporal coherence from novel viewpoints. ReconDreamer [ni2025recondreamer] introduces an online restoration module together with progressive data reuse to correct artifacts in Gaussian-rendered views and enables reliable reconstruction of large-scale trajectories. MagicDrive3D [gao2024magicdrive3d] generates multi-view street scenes conditioned on BEV maps, 3D boxes, and text, and further converts the outputs into full 3D environments through fault-tolerant GS. In contrast, implicit-field methods replace GS with continuous neural representations. UnO [agro2024uno] leverages future point clouds to learn a NeRF-style 4D occupancy field, which allows annotation-free prediction and achieves strong transfer performance beyond supervised baselines in point-cloud forecasting.
IV Data Resources & Metrics
World models in embodied AI are required to address diverse tasks spanning manipulation, navigation, and autonomous driving, requiring heterogeneous resources and rigorous evaluation. Accordingly, we present Data Resources in §IV-A and Metrics in §IV-B, focusing on the most widely adopted platforms and evaluation measures as a unified foundation for cross-domain assessment.
IV-A Data Resources